OpenStudy (anonymous):

specify the domain for the function f(x)=1/x^2-5x+25 a.)(-5/2,5/2) b.)(-infinity,infinity) c.)(-infinity,5-square root of 10/2)U(5+square root of 10/2, infinity) d.)(-infinity, 5-5 square root of 3/2)U(5+5 square root of 3/2, infinity

OpenStudy (anonymous):

yikes

OpenStudy (anonymous):

set \[x^2-5x+25=0\] and solve for x

OpenStudy (anonymous):

since there is no real solution to that equation, go with \((-\infty, \infty)\)

OpenStudy (anonymous):

thank u!